# A uniform-in-time nonlocal approximation of the standard Fokker-Planck equation

### v2

@@ Abstract
We study a nonlocal approximation of the Fokker-Planck equation in which we can estimate the speed of convergence to equilibrium in a way which respects the local limit of the equation. This uniform estimate cannot be easily obtained with standard inequalities or entropy methods, but can be obtained through the use of Harris’s theorem, finding interesting links to quantitative versions of the central limit theorem in probability. As a consequence one can prove that solutions of this nonlocal approximation converge to solutions of the usual Fokker-Planck equation uniformly in time—hence we show the approximation is asymptotic-preserving in this sense. The associated equilibrium has some interesting tail and regularity properties, which we also study.
@@ 1 Introduction
for $J$ a general probability distribution satisfying (1). We have chosen to write the PDE as in (NLFP) in order to empasise that we are especially interested in its scaling limit to the Fokker-Planck equation.
We are not aware of any studies regarding the specific PDE (NLFP) except for Mischler and Tristani (2017), whose results are similar to our Theorem 1.1, but use very different techniques (our conditions on $J$ are somewhat simplified, allowing less regular and non-symmetric $J$). Our result in Theorem 1.2 regarding the nonlocal-to-local limit as $\varepsilon\to 0$ is new as far as we know. There are plenty of results on related equations, of which we give a short summary now, but we highlight that nonlocal-to-local limits which preserve the asymptotic behavior in the limit are comparatively rare, and are a current source of interest in several contexts (this is also the point of view of Mischler and Tristani (2017), where other nonlocal PDE are investigated as well).
The approach from nonlocal diffusion-type operators of the form $J*u-u$ to local diffusion operators has been the subject of many previous works. The PDE $\partial_{t}u=J*u-u$ is studied in Chasseigne et al. (2006), and a review of related techniques (making strong use of explicit solutions via the Fourier transform) is given in Andreu-Vaillo et al. (2010). Generalizations of this result to similar nonlocal PDE with different integral kernels are given in Molino and Rossi (2019). The same PDE is studied in Rey and Toscani (2013), with a focus on its large-time behavior. In this case the motivation comes from both physics (since this PDE is thought to be a more appropriate diffusion model in some cases) and numerical analysis (see below for more on this). Some results on its large-time behavior can be obtained by using Fourier-type distances (Carrillo and Toscani, 2007), again by taking advantage of the fact that the PDE is explicitly solvable in Fourier variables. A different nonlocal and nonlinear version of the Fokker-Planck PDE (FP) is studied in Ignat and Rossi (2007). Its relationship to the local problem is also studied there, but the results on asymptotic behavior are not preserved when passing to the local limit. Still a different nonlocal version has been recently considered in Auricchio et al. (2023), motivated by a model in collective behavior.
Let us describe more precisely the results we are able to show. We denote by $B_{r}(x_{0})$, the ball in $\mathbb{R}^{d}$ centered at $x_{0}$ with radius $r$. When the ball is centered at 00, we use $B_{r}$ as a shorthand. We always denote by $L_{1}^{k}$ the usual weighted Lebesgue space with norm
@@
where $\left\langle x\right\rangle^{k}:=(1+|x|^{2})^{k/2}$, a “smoothed $k$-th power” which avoids regularity problems at $x=0$. We denote by $\left\langle x\right\rangle$ the smoothed $1^{st}$ power. Similarly we define $L^{1}_{\exp,a}$ as the Lebesgue space with weight $e^{a|x|}$ with natural norm
@@
$\frac{1}{\varepsilon^{2}}\big{(}J_{\varepsilon}*F-F\big{)}+\operatorname{div}(xF)=0$
@@
In fact, from our proof of the theorem (or directly as a simple consequence) we get the same estimates substituting $u-F_{\varepsilon}$ by any solution $w$ with initial mass 00 (i.e. $\int_{\mathbb{R}^{d}}w_{0}(x)\,\mathrm{d}x=0$). For example, we prove that under the conditions of point (i),
@@
Proving an estimate like the previous one has several consequences. One of these is the following uniform-in-time result on the closeness solutions to the local and nonlocal problems, which we prove in Section 4. For readability, we state the theorem with likely non-optimal hypotheses, since optimising it is not the main goal of this paper. It requires the following additional hypothesis on $J$:
$\int_{\mathbb{R}^{d}}J(x)x^{\alpha}=0\qquad\text{for all multiindices $\alpha$of order $|\alpha|=3$.}$ (4)
(Which is satisfied for example if $J$ is symmetric.)
@@
$\left\|u(\cdot,t)-v(\cdot,t)\right\|_{L^{1}_{k}}\leq C\varepsilon^{2}\qquad\text{for every $t\geq 0$ and $\varepsilon\in(0,1]$.}$ (5)
@@
$\left\|u(\cdot,t)-v(\cdot,t)\right\|_{L^{1}_{k}}\leq C\varepsilon\qquad\text{for every $t\geq 0$ and $\varepsilon\in(0,1]$,}$ (6)
for some (other) constant $C>0$ depending only on $d$ and $u_{0}$. In order to obtain this it is enough to carry out the Taylor expansions in Section 4 to one order less.
@@
We empasise once more that, although being far from optimal, the constants we find are all explicit, since proofs are completely constructive.
We will also make a few remarks regarding the unique probability equilibrium $F_{\varepsilon}$ associated to (NLFP). Theorems 1.1 and 1.2 imply in particular that $\|F_{\varepsilon}-G\|_{L^{1}_{k}}\leq C\varepsilon$, where $G$ is the standard Gaussian probability distribution in $\mathbb{R}^{d}$. But the properties of $F_{\varepsilon}$ are quite interesting in their own right. Theorem 1.1 includes the fact that $F_{\varepsilon}$ has exponentially-decaying tails as long as $J$ has the same property. However, when $J$ is compactly supported we suspect that $F_{\varepsilon}$ has tails which decay like the Poisson distribution, that is $F_{\varepsilon}\approx e^{-a|x|\log|x|}$ for large $|x|$. We are able to show an incomplete version of this statement in Section 5.3. We suspect that in general $F_{\varepsilon}$ does not have Gaussian tails, but we haven’t been able to prove this.
The regularity of $F_{\varepsilon}$ is also interesting, as we explore in Section 5.1: independently of the regularity of $J$, $F_{\varepsilon}$ appears to be smooth everywhere except at $x=0$; and its regularity at $x=0$ depends on $\varepsilon$, not on the regularity or tails of $J$. The equilibrium $F_{\varepsilon}$ becomes more regular as $\varepsilon\to 0$. We think a more complete study of the properties of $F_{\varepsilon}$ would be interesting for a future work.
@@
$u(t)(x)=\sum_{n=0}^{\infty}\Big{(}\frac{1}{\varepsilon^{2}}\Big{)}^{n}\int_{0}^{t}\int_{0}^{t_{1}}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}(e^{t}x)\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1},$
@@ 2.1 Existence and Uniqueness
Let $k\geq 0$, and take $J$ satisfying Hypothesis (1) on $J$ and $J\in L^{1}_{k}$. Take $u_{0}\in L^{1}_{k}$, and take $T^{*}\in(0,+\infty]$. A mild solution of equation (9) on $[0,T^{*})$ with initial condition $u_{0}$ is a function $u\in\mathcal{C}([0,T^{*}),L^{1}_{k})$ such that
@@ 2.1.1 Existence of $L^{1}_{k}$ solutions
For every $T^{*}$, there exists $n\in\mathbb{N}$ such that $\big{(}\frac{c_{k}T^{*}}{\varepsilon^{2}}\big{)}^{n}\frac{1}{n!}<1$. There is a specific version of the contraction mapping theorem (see for example Miklavcic (1998, Thm. 1.1.3)) which now implies $\Psi$ has a fixed point which is a solution of (NLFP) in the sense of Definition 2.1. (Alternatively, one can consider the sequence $(\Psi^{n}(u_{0}))_{n\geq 0}$, where $u_{0}\in\mathcal{Y}$ is the function constantly equal to $u_{0}$ on $[0,T^{*}]$; our last inequality easily shows that it is a Cauchy sequence, and hence must have a limit. This can be done for any $T^{*}>0$, easily obtaining the result on $[0,+\infty)$.)
@@
By the standard Gronwall inequality applied to $e^{\frac{t}{\varepsilon^{2}}}\left\|u(t)-v(t)\right\|_{L^{1}_{k}}$ we obtain the result. ∎
@@ 2.1.2 Existence of measure solutions
We sketch similar arguments that give existence of solutions in a space of measures, since the Harris theorem we use later is usually stated for semigroups in a space of measures (or Markov semigroups in the context of probability). This is really not strictly necessary, since there are versions of the Harris theorem which hold similarly in weighted $L^{1}$ spaces (see Cañizo and Mischler (2023)), but since the proof is very similar we give the main steps.
@@
where :=$\mathcal{F}:=\{f\in W^{1,\infty}(\mathbb{R}^{d}):\left\|f\right\|_{1,\infty}\leq 1\}$, with $\left\|f\right\|_{1,\infty}:=\|f\|_{\infty}+\operatorname{Lip}(f)$ and $\operatorname{Lip}(f)$ denoting the Lipschitz constant of $f$. In the literature, the bounded Lipschitz norm is sometimes defined as the equivalent norm obtained by choosing in the previous definition :=$\mathcal{F}:=\{f:\mathbb{R}^{d}\to\mathbb{R}:\left\|f\right\|_{\infty}\leq 1,\sup_{x\neq y}\frac{\left|f(x)-f(y)\right|}{|x-y|}\leq 1\}$.
We define the ball $\mathcal{X}_{R}:=\{\mu\in\mathcal{M}:\left\|\mu\right\|_{\mathcal{M}_{k}}\leq R\}$ and the set $\mathcal{Z}=\mathcal{C}([0,T^{*}],\mathcal{X}_{R})$ of functions which are continuous from $[0,T^{*}]$ to $\mathcal{X}_{R}$ in the bounded Lipschitz norm. $\mathcal{Z}$ is a complete metric space with the metric induced by the norm
@@
We have left to show that $\Psi[\mu](t)\in\mathcal{X}_{R}$ for every $t\in[0,T^{*}]$ when $\mu(t)\in\mathcal{X}_{R}$. Set $R=2\left\|u_{0}\right\|_{\mathcal{M}_{k}}$
@@ 2.2 Semigroup results
Our result on the well-posedness of the nonlocal Fokker-Planck equation (NLFP) can also be phrased in terms of semigroup theory. For an introduction and further details on the results we will be using below, we refer the reader to the books Engel and Nagel (2001) or Pazy (1983). We first point out that for any $k\geq 0$, the operators $(T_{t})_{t\geq 0}$ given in (7) define a strongly continuous semigroup on the Banach space $L^{1}_{k}$; it is easy to check this directly. The generator of this semigroup is the operator $B$ given by
@@ 2.3 Representation of the solution
with the initial condition $F_{\varepsilon}(0)=1$, since $\left\|F_{\varepsilon}\right\|_{L^{1}}=1$. In particular
@@
$h_{n}(t)=h_{n-1}(t)+\frac{1}{\varepsilon^{2n}}\int_{0}^{t}\int_{0}^{t_{1}}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1},$
where we use as a shorthand the following
$J_{\varepsilon}^{t_{1},\dots,t_{n}}:=J_{\varepsilon e^{t_{1}}}*J_{\varepsilon e^{t_{2}}}*\cdots*J_{\varepsilon e^{t_{n}}}$
@@
From the previous is immediate to see that if $u_{0}$ and $J$ are radially symmetric functions then also the solution is radially symmetric. One can also say this of any property which is preserved by convolution with $J$. Analogously, the equilibrium will preserve the same properties.

### v3

@@ Abstract
We study a nonlocal approximation of the Fokker-Planck equation in which we can estimate the speed of convergence to equilibrium in a way which does not degenerate as we approach the local limit of the equation. This uniform estimate cannot be easily obtained with standard inequalities or entropy methods, but can be obtained through the use of Harris’s theorem, finding interesting links to quantitative versions of the central limit theorem in probability. As a consequence one can prove that solutions of this nonlocal approximation converge to solutions of the usual Fokker-Planck equation uniformly in time—hence we show the approximation is asymptotic-preserving in this sense. The associated equilibrium has some interesting tail and regularity properties, which we also study.
@@ 1 Introduction
for $J$ a general probability distribution satisfying (1). We have chosen to write the PDE as in (NLFP) in order to emphasize that we are especially interested in its scaling limit to the Fokker-Planck equation.
Summary of main results
Here are the two main results we will prove in this paper:
  1. 1.
    Solutions to (NLFP) converge to equilibrium exponentially:
    $\|u(t,\cdot)-F_{\varepsilon}\|\leq Ce^{-\gamma t}\|u_{0}-F_{\varepsilon}\|,\qquad t\geq 0$
    in a variety of weighted $L^{1}$ norms $\|\cdot\|$, where $F_{\varepsilon}$ is a suitable equilibrium of the PDE. Our main point is that both constants $C$ and $\gamma$ can be taken to be independent of $\varepsilon$ as $\varepsilon\to 0$. This is not contained in previous results, except for Mischler and Tristani (2017), who use different techniques.
  2. 2.
    If $u$ is the solution to the nonlocal equation (NLFP) with a symmetric $J$ and $v$ the solution to the standard Fokker-Planck equation (FP) with the same initial condition, then
    $\|u(t,\cdot)-v(t,\cdot)\|\leq C\varepsilon^{2}\qquad\text{for all $t\geq 0$,}$
    again in a variety of weighted $L^{1}$ norms. The constant $C$ is independent of $\varepsilon$ and $t$, so the convergence happens uniformly for all times. We refer to this as “asymptotic-preserving convergence”, and is not proved in previous literature as far as we know.
Let us describe these results more precisely. We always denote by $L_{1}^{k}$ the usual weighted Lebesgue space with norm
@@
where $\left\langle x\right\rangle^{k}:=(1+|x|^{2})^{k/2}$, a “smoothed $k$-th power” which avoids regularity problems at $x=0$. (We denote by $\left\langle x\right\rangle$ the smoothed $1^{st}$ power.) Similarly we define $L^{1}_{\exp,a}$ as the Lebesgue space with weight $e^{a|x|}$ with natural norm
@@
$\frac{1}{\varepsilon^{2}}\big(J_{\varepsilon}*F_{\varepsilon}-F_{\varepsilon}\big)+\operatorname{div}(xF_{\varepsilon})=0$
@@
The dependence of the constants can be explicitly traced throughout the proof. Among the various factors involved, the constants in general get better the closer $J$ resembles a Gaussian, in a sense that will be clarified later.
Remark 1.2.
From our proof of the theorem (or directly as a simple consequence) we get the same estimates substituting $u-F_{\varepsilon}$ by any solution $w$ with initial mass 0 (i.e. $\int_{\mathbb{R}^{d}}w_{0}(x)\,\mathrm{d}x=0$). For example, we prove that under the conditions of point (i),
@@
Proving an estimate like the previous one has several consequences. One of these is the following uniform-in-time result on the closeness solutions to the local and nonlocal problems, which we prove in Section 4. For readability, we state the theorem with likely non-optimal hypotheses, since optimizing it is not the main goal of this paper. It requires the following additional hypothesis on $J$:
$\int_{\mathbb{R}^{d}}J(x)x^{\alpha}\,\mathrm{d}x=0\qquad\text{for all multiindices $\alpha$ of order $|\alpha|=3$,}$ (4)
which is satisfied for example if $J$ is symmetric.
@@
$\left\|u(t,\cdot)-v(t,\cdot)\right\|_{L^{1}_{k}}\leq C\varepsilon^{2}\qquad\text{for every $t\geq 0$ and $\varepsilon\in(0,1]$.}$ (5)
@@
$\left\|u(t,\cdot)-v(t,\cdot)\right\|_{L^{1}_{k}}\leq C\varepsilon\qquad\text{for every $t\geq 0$ and $\varepsilon\in(0,1]$,}$ (6)
for some (other) $\varepsilon$-independent constant $C>0$. In order to obtain this it is enough to carry out the Taylor expansions in Section 4 to one order less.
@@
We emphasize once more that, although being far from optimal, the constants we find are all explicit, since proofs are completely constructive.
We will also make a few remarks regarding the unique probability equilibrium $F_{\varepsilon}$ associated to (NLFP). Theorems 1.1 and 1.3 imply in particular that $\|F_{\varepsilon}-G\|_{L^{1}_{k}}\leq C\varepsilon^{2}$, where $G$ is the standard Gaussian probability distribution in $\mathbb{R}^{d}$. But the properties of $F_{\varepsilon}$ are quite interesting in their own right. Theorem 1.1 includes the fact that $F_{\varepsilon}$ has exponentially-decaying tails as long as $J$ has the same property. However, when $J$ is compactly supported we suspect that $F_{\varepsilon}$ has tails which decay like the Poisson distribution, that is $F_{\varepsilon}\approx e^{-a|x|\log|x|}$ for large $|x|$. We are able to show an incomplete version of this statement in Section 5.3. We suspect that in general $F_{\varepsilon}$ does not have Gaussian tails, but we haven’t been able to prove this.
The regularity of $F_{\varepsilon}$ is also interesting, as we explore in Section 5.1: independently of the regularity of $J$, $F_{\varepsilon}$ appears to be smooth everywhere except at $x=0$; and its regularity at $x=0$ depends on $\varepsilon$, not on the regularity nor tails of $J$. The equilibrium $F_{\varepsilon}$ becomes more regular as $\varepsilon\to 0$. We think a more complete study of the properties of $F_{\varepsilon}$ would be interesting for a future work.
@@
$u(t)(x)=e^{(d-\frac{1}{\varepsilon^{2}})t}\sum_{n=0}^{\infty}\Big(\frac{1}{\varepsilon^{2}}\Big)^{n}\int_{0}^{t}\int_{0}^{t_{1}}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}(e^{t}x)\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1},$
@@
Previous literature and related results
We are not aware of any studies regarding the specific PDE (NLFP) except for Mischler and Tristani (2017), whose results are similar to our Theorem 1.1, but use very different techniques (our conditions on $J$ are somewhat simplified, allowing less regular and non-symmetric $J$). Our result in Theorem 1.3 regarding the nonlocal-to-local limit as $\varepsilon\to 0$ is new as far as we know. There are plenty of results on related equations, of which we give a short summary now, but we highlight that nonlocal-to-local limits which preserve the asymptotic behavior in the limit are comparatively rare, and are a current source of interest in several contexts.
The approach from nonlocal diffusion-type operators of the form JuuJ*u-u to local diffusion operators has been the subject of many previous works. The PDE $\partial_{t}u=J*u-u$ is studied in Chasseigne et al. (2006), and a review of related techniques (making strong use of explicit solutions via the Fourier transform) is given in Andreu-Vaillo et al. (2010). Generalizations of this result to similar nonlocal PDE with different integral kernels are given in Molino and Rossi (2019). The same PDE is studied in Rey and Toscani (2013), with a focus on its large-time behavior. In this case the motivation comes from both physics (since this PDE is thought to be a more appropriate diffusion model in some cases) and numerical analysis (see below for more on this). Some results on its large-time behavior can be obtained by using Fourier-type distances (Carrillo and Toscani, 2007), again by taking advantage of the fact that the PDE is explicitly solvable in Fourier variables. A different nonlocal and nonlinear version of the Fokker-Planck PDE (FP) is studied in Ignat and Rossi (2007). Its relationship to the local problem is also studied there, but the results on asymptotic behavior are not preserved when passing to the local limit. Still a different nonlocal version has been recently considered in Auricchio et al. (2023), motivated by a model in collective behavior.
@@ 2.1 Existence and Uniqueness
Let $k\geq 0$, and take $J$ satisfying Hypothesis (1) on $J$ and $J\in L^{1}_{k}$. Take $u_{0}\in L^{1}_{k}$, and take $T^{*}\in(0,+\infty]$. A mild solution of equation (NLFP) on $0,T^{*})$ with initial condition $u_{0}$ is a function $u\in\mathcal{C}([0,T^{*}),L^{1}_{k})$ such that
@@ 2.1.1 Existence of $L^{1}_{k}$ solutions
For every $T^{*}$, there exists $n\in\mathbb{N}$ such that $\big(\frac{c_{k}T^{*}}{\varepsilon^{2}}\big)^{n}\frac{1}{n!}<1$. There is a specific version of the contraction mapping theorem (see for example Miklavcic (1998, Thm. 1.1.3)) which now implies $\Psi$ has a fixed point which is a solution of (NLFP) in the sense of Definition 2.1. (Alternatively, one can consider the sequence $(\Psi^{n}(u_{0}))_{n\geq 0}$, where $u_{0}\in\mathcal{Y}$ is the function constantly equal to $u_{0}$ on $0,T^{*}]$; our last inequality easily shows that it is a Cauchy sequence, and hence must have a limit. This can be done for any $T^{*}>0$, easily obtaining the result on $0,+\infty)$.)
@@
By the standard Gronwall lemma applied to $e^{\frac{t}{\varepsilon^{2}}}\left\|u(t)-v(t)\right\|_{L^{1}_{k}}$ we obtain the result. ∎
@@ 2.1.2 Existence of measure solutions
We sketch similar arguments that give existence of solutions in a space of measures, since Harris’s theorem we use later is usually stated for semigroups in a space of measures (or Markov semigroups in the context of probability). This is really not strictly necessary, since there are versions of the Harris theorem which hold similarly in weighted $L^{1}$ spaces (see Cañizo and Mischler (2023)), but since the proof is very similar we give the main steps.
@@
where $\mathcal{F}:=\{f\in W^{1,\infty}(\mathbb{R}^{d}):\left\|f\right\|_{1,\infty}\leq 1\}$, with $\left\|f\right\|_{1,\infty}:=\|f\|_{\infty}+\operatorname{Lip}(f)$ and $\operatorname{Lip}(f)$ denoting the Lipschitz constant of $f$. In the literature, the bounded Lipschitz norm is sometimes defined as the equivalent norm obtained by choosing in the previous definition $\mathcal{F}:=\{f:\mathbb{R}^{d}\to\mathbb{R}:\left\|f\right\|_{\infty}\leq 1,\sup_{x\neq y}\frac{\left|f(x)-f(y)\right|}{|x-y|}\leq 1\}$.
We define the ball $\mathcal{X}_{R}:=\{\mu\in\mathcal{M}:\left\|\mu\right\|_{\mathcal{M}_{k}}\leq R\}$ and the set $\mathcal{Z}=\mathcal{C}([0,T^{*}],\mathcal{X}_{R})$ of functions which are continuous from $0,T^{*}]$ to $\mathcal{X}_{R}$ in the bounded Lipschitz norm. $\mathcal{Z}$ is a complete metric space with the metric induced by the norm
@@
We have left to show that $\Psi[\mu](t)\in\mathcal{X}_{R}$ for every $t\in[0,T^{*}]$ when $\mu(t)\in\mathcal{X}_{R}$. Set $R=2\left\|\mu_{0}\right\|_{\mathcal{M}_{k}}$
@@ 2.2 Semigroup results
Our result on the well-posedness of the nonlocal Fokker-Planck equation (NLFP) can also be phrased in terms of semigroup theory. For an introduction and further details on the results we will be using below, we refer the reader to the books Engel and Nagel (2001) or Pazy (1983). We first point out that for any $k\geq 0$, the operators $(T_{t})_{t\geq 0}$ given in (7) define a strongly continuous semigroup on the Banach space $L^{1}_{k}$; it is easy to check this directly. The generator of this semigroup is the operator $B_{\varepsilon}$ given by
@@ 2.3 Representation of the solution
with the initial condition $\hat{F}_{\varepsilon}(0)=1$, since $\left\|F_{\varepsilon}\right\|_{L^{1}}=1$. In particular
@@
$h_{n}(t)=h_{n-1}(t)+\frac{1}{\varepsilon^{2n}}\int_{0}^{t}\int_{0}^{t_{1}}\dots\int_{0}^{t_{n-1}}J_{\varepsilon}^{t_{1},\dots,t_{n}}*u_{0}\,\mathrm{d}t_{n}\dots\,\mathrm{d}t_{1}.$
@@
From the previous it is immediate to see that if $u_{0}$ and $J$ are radially symmetric functions then also the solution is radially symmetric. One can also say this of any property which is preserved by convolution with $J$. Analogously, the equilibrium will preserve the same properties.